An electric vehicle starts from rest and accelerates at a rate of 1.8 m/s2 in a straight line until it reaches a speed of 18 m/s. The vehicle then slows at a constant rate of 1.4 m/s2 until it stops. (a) How much time elapses from start to stop? (b) How far does the vehicle move from start to stop?
Answer #1

a)

During acceleration :

vo = initial velocity = 0 m/s

a = acceleration = 1.8 m/s2

v = final velocity = 18 m/s

t = time taken

Using the equation

v = vo + a t

18 = 0 + (1.8) t

t = 10 sec

During retardation :

v'o = initial velocity = 18 m/s

a' = acceleration = - 1.4 m/s2

v' = final velocity = 0 m/s

t' = time taken

Using the equation

v' = v'o + a' t'

0 = 18 + (- 1.4) t'

t' = 12.86  sec

T = Total time = t + t' = 10 + 12.86 = 22.86 sec

b)

During acceleration :

vo = initial velocity = 0 m/s

a = acceleration = 1.8 m/s2

v = final velocity = 18 m/s

d = distance traveled  

Using the equation

v2 = v2o + 2 a d

182 = 02 + 2 (1.8) d

d = 90 m

During retardation :

v'o = initial velocity = 18 m/s

a' = acceleration = - 1.4 m/s2

v' = final velocity = 0 m/s

d' = distance traveled

Using the equation

v'2 = v'2o + 2 a' d'

02 = 182 + 2 (- 1.4) d'

d' = 115.71 m

D = Total distance traveled = d + d' = 90 + 115.7 = 205.7 m